# LongCaptioning: Unlocking the Power of Long Caption Generation in Large Multimodal Models

### v1

@@ 4 LongCaption-Bench: A Benchmark for Long-Caption Generation
To reliably evaluate the long-caption generation capabilities of LMMs, we constructed a LongCaption-Bench. Specifically, to prevent test videos from being included in the training data, we selected videos from the test sets of open-source long video benchmarks. To reduce the difficulty of manual annotation, we first used the latest proprietary LMM, Gemini 1.5 Pro [34] (which supports long-caption generation), to generate an initial caption for all videos. We then manually reviewed each video and its corresponding initial caption, removing any abnormal captions and their associated samples. Abnormal captions included those with infinite repetition, incomplete captions (abrupt termination), and sensitive content. In the end, we retained 281 valid video-caption pairs. Subsequently, we manually reviewed the initial captions against the video content and corrected or supplemented any errors or omissions in the captions. Table 2 reports the statistics of LongCaption-Bench. The test videos are primarily concentrated in the 900 to 1,200-second range, with caption lengths mostly between 500 and 1,500 words. The average video duration is 1060.4s, and the average caption length is 1161.3 words, which is significantly greater than previous video captioning benchmarks.
@@
Table 3: Evaluation results on the LongCaption-Bench across different duration intervals. Here, $S_{l}$ represents the score for caption length, $S_{q}$ represents the quality score, and $S_{r}$ represents the video-caption relevance score.
Overall [300s, 600s) [600s, 900s) [900s, 1200s) [1200s, 1800s)
$S_{l}$ $S_{q}$ $S_{r}$ $S_{l}$ $S_{q}$ $S_{r}$ $S_{l}$ $S_{q}$ $S_{r}$ $S_{l}$ $S_{q}$ $S_{r}$ $S_{l}$ $S_{q}$ $S_{r}$
Proprietary models
GPT-4o [32] 33.6 79.4 2.50 64.9 81.5 2.78 41.2 81.2 2.41 30.2 79.4 2.57 24.3 78.7 2.50
Open-source models
PLLaVA 7B [41] 4.5 63.2 1.10 4.0 64.1 1.09 2.3 63.7 1.07 3.7 64.9 1.10 6.8 62.8 1.18
PLLaVA 32B [41] 5.3 65.1 1.18 4.9 66.4 1.21 2.6 64.7 1.10 4.9 65.2 1.20 7.7 64.0 1.17
LongVA 7B [47] 2.1 64.5 1.15 0.0 65.2 1.11 0.0 65.5 1.07 1.9 64.6 1.17 4.0 63.6 1.18
LLaVA-Video 7B [27] 4.5 66.8 1.30 8.6 67.8 1.33 1.8 69.0 1.21 3.1 67.2 1.38 8.4 64.8 1.24
MiniCPMV2.6 8B [45] 11.8 73.5 2.18 27.6 72.6 2.11 12.8 74.9 2.02 10.4 73.1 2.24 11.5 73.0 2.20
Qwen2-VL 7B [37] 7.7 57.8 1.36 0.0 58.9 1.22 6.6 56.4 1.12 6.4 58.5 1.47 11.2 57.6 1.36
Qwen2-VL 72B [37] 9.4 70.2 1.71 11.8 69.1 1.72 7.7 68.6 1.66 7.0 71.4 1.72 10.1 70.2 1.68
Our trained models
LongCaptioning-8B 40.9 81.2 2.59 88.2 80.7 2.56 53.6 81.7 2.53 38.8 81.6 2.62 30.3 80.4 2.58
@@
Table 4: Ablation study results on the LongCaption-Bench. ‘w/o Visual Context Window Ext.’ represents the LongCaptioning 8B model trained without the visual context window extension, ‘w/o Clip-Level Captioning’ indicates that only frame-level captions were used during long caption synthesis, ‘w/o Frame-Level Captioning’ refers to the use of only clip-level captions during long caption synthesis, and ‘w/o LongCaption-10k data’ refers to the results without using the LongCaption-10K dataset for training.
Overall [300s, 600s) [600s, 900s) [900s, 1200s) [1200s, 1800s)
$S_{l}$ $S_{q}$ $S_{r}$ $S_{l}$ $S_{q}$ $S_{r}$ $S_{l}$ $S_{q}$ $S_{r}$ $S_{l}$ $S_{q}$ $S_{r}$ $S_{l}$ $S_{q}$ $S_{r}$
LongCaptioning-8B 40.9 81.2 2.59 88.2 80.7 2.56 53.6 81.7 2.53 38.8 81.6 2.62 30.3 80.4 2.58
   w/o Visual Context Window Ext. 32.3 79.1 2.35 90.3 76.3 2.44 37.5 78.7 2.13 30.1 79.7 2.37 25.8 78.6 2.45
   w/o Clip-Level Captioning 28.7 75.0 2.21 64.3 74.8 2.25 25.8 76.5 2.06 21.0 76.7 2.28 19.1 74.9 2.30
   w/o Frame-Level Captioning 25.7 74.2 2.20 58.8 74.6 2.20 21.5 76.1 2.06 18.9 76.0 2.27 16.0 74.3 2.28
   w/o LongCaption-10k data 11.8 73.5 2.18 27.6 72.6 2.11 12.8 74.9 2.02 10.4 73.1 2.24 11.5 73.0 2.20
5 LongCaptioning: A Long-Caption Generation Model
@@ 5.1 Model Training
We conduct training based on the latest open-source models, namely MiniCPMV2.6 8B [45]. This model has only undergone pretraining and supervised fine-tuning (SFT) on image-text datasets. This allows us to eliminate the interference of other video datasets when analyzing the experimental results. Moreover, to further extend the effective context window during inference and allow the input of longer frame sequences, we introduced a visual context window extension technique [39] during training. This technique scales the rotational frequency of the position encoding embedding for visual tokens, thereby extending the model’s context window. The model was trained on a 2 A800-80G GPUs using DeepSpeed + ZeRO3 + CPU offloading [2]. We used a batch size of 1, a learning rate of 5e-5, and trained for 1 epoch.
@@ 5.2.1 Quatitative Results
Evaluates the Output Length Table 3 presents the results of the output length score of the models. For the output length score $S_{l}$, we use human annotations as the ground truth (assuming the human annotations contain all the content of the test videos) and calculate the difference between the model output length and the human annotation length using Equation 2. A higher $S_{l}$ indicates that the length of the captions generated by the model more closely matches the length of the human annotations. Our findings are as follows: 1) Open-source LMMs generally perform poorly, even those with 72 billion parameters, as they struggle to generate complete captions beyond 300 words. 2) The performance of MiniCPMV2.6 8B even surpasses that of Qwen2-VL 72B, reflecting that the output length of LMMs is not necessarily correlated with model size. 3) Notably, compared to other duration intervals, all models show an improvement in output length scores in the [1200s, 1800s) interval. After analyzing the generated outputs, we found that some models exhibit repetitive output issues when the input frame is longer, leading to an increase in output length. Due to the repetition problem, their quality scores are lower. 4) By fine-tuning the model on LongCaption-10K, our model achieved optimal performance, even surpassing larger proprietary models like GPT-4o.
Figure 5 presents the output length statistics for various models, with anomalous outputs like infinitely looping captions manually removed. Open-source models, such as MiniCPMV2.6 8B and Qwen2-VL 72B, generated captions primarily between 150 and 300 words. The proprietary GPT-4o model produced captions mostly in the 300 to 600-word range. In contrast, our model generated outputs between 300 and 700 words, with a maximum length reaching approximately 1,200 words. This demonstrates our model’s ability to produce longer, more detailed captions compared to both open-source and proprietary models.
@@
Evaluates the Relevance of Video and Generated Caption To assess the relevance of model-generated descriptions to the video content, Table 3 presents relevance scores $S_{r}$. Our model consistently achieved the highest relevance scores across all duration intervals, indicating its superior ability to accurately and comprehensively describe video content compared to other models.
@@ 5.2.2 Ablation Studies
We explore the effects of removing the visual context window extension and the LongCaption-10K dataset on the LongCaptioning-8B model. As shown in Table 4, we find that: 1) Removing the visual context window extension significantly reduces both output length and quality, with the length score $S_{l}$ dropping by 8.6, the quality score $S_{q}$ dropping by 2.1, and the relevance score $S_{r}$ dropping by 0.24. This is because the visual context window extension allows the model to input more sampled frames during the inference process. 2) Removing the clip-level captioning step during the synthesis of LongCaption-10K leads to a drop in all metrics, as relying solely on frame-level captions fails to capture dynamic changes and action continuity, resulting in less coherent captions. 3) Similarly, removing the frame-level captioning step further degrades performance, as using only clip-level captions during LongCaption-10K synthesis results in shorter captions. 4) Without the LongCaption-10K dataset ($i.e.$, using only the MiniCPMV2.6 8B backbone), the model struggles to generate long captions, leading to a sharp decline in output length. Additionally, without both the visual context window extension and the LongCaption-10K dataset, the relevance between generated captions and video content significantly deteriorates.
5.2.3 Qualitative Results
Figure 6 illustrates the qualitative results of video captions generated by the comparison models (MiniCPMV2.6 8B and Qwen2-VL 72B) and our model (LongCaptioning 8B). For all models, we uniformly sample 128 frames from the video as input. It is evident that LongCaptioning 8B produces outputs with significantly greater length compared to the other models while maintaining the relevance of the generated content to the video. In contrast to MiniCPMV2.6 8B and Qwen2-VL 72B, LongCaptioning 8B generates more detailed captions for the input video.
@@ 6 Conclusions
Supplementary Material
7 Training Details
@@
Table 5 reports the details of the training process. We conduct training based on the latest open-source models, namely MiniCPMV2.6 8B [45]. This model has only undergone pretraining and supervised fine-tuning (SFT) on image-text datasets. This allows us to eliminate the interference of other video datasets when analyzing the experimental results. We employ the LoRA strategy for supervised fine-tuning. Specifically, we fine-tune only a small subset of parameters in the LLM, where the $\text{LoRA}_{r}$ is set to 8 and the $\text{LoRA}_{\alpha}$ is set to 32. Moreover, to further extend the effective context window during inference and allow the input of longer frame sequences, we introduced a visual context window extension technique [39] during training. This technique scales the rotational frequency of the position encoding embedding for visual tokens, thereby extending the model’s context window. The model was trained on a 2 A800-80G GPUs using DeepSpeed + ZeRO3 + CPU offloading [2]. We used a batch size of 1, a learning rate of 5e-5, and trained for 1 epoch.
8 Prompts
8.1 Test Prompt
In Figure 2 and 2, all models use the same prompt.
Task: You are an expert in understanding scene transitions based on visual features in a video. You are requested to create the descriptions for the current video sent to you, which includes multiple sequential frames. Guidelines For Video Description: - Analyze the narrative progression implied by the sequence of frames, interpreting the sequence as a whole. - Note that since these frames are extracted from a video, adjacent frames may show minimal differences. These should not be interpreted as special effects in the video. - If text appears in the frames, you must describe the text in its original language and provide an English translation in parentheses. For example: book. Additionally, explain the meaning of the text within its context. - When referring to people, use their characteristics, such as clothing, to distinguish different people. - **IMPORTANT** Please provide as many details as possible in your description, including colors, shapes, and textures of objects, actions and characteristics of humans, as well as scenes and backgrounds. Output Format: Your response should look like this: ”Video Level Description”: ”The video begins with…, progresses by…, and concludes with…” Please give me the description of the current video.
8.2 Quality Score
@@
You are an expert in evaluating text quality. Please evaluate the quality of an AI assistant’s response to a user’s writing request. Be as strict as possible. You need to evaluate across the following six dimensions, with scores ranging from 1 to 5. The scoring criteria from 5 to 1 for each dimension are as follows: 1. Relevance: From content highly relevant and fully applicable to the user’s request to completely irrelevant or inapplicable. 2. Accuracy: From content completely accurate with no factual errors or misleading information to content with numerous errors and highly misleading. 3. Coherence: From clear structure with smooth logical connections to disorganized structure with no coherence. 4. Clarity: From clear language, rich in detail, and easy to understand to confusing expression with minimal details. 5. Breadth and Depth: From both broad and deep content with a lot of information to seriously lacking breadth and depth with minimal information. 6. Reading Experience: From excellent reading experience, engaging and easy to understand content to very poor reading experience, boring and hard to understand content. Please evaluate the quality of the following response to a user’s request according to the above requirements. $\langle User\quad Request\rangle$ : $\langle/Response\rangle$ : ”’Please evaluate the quality of the response. You must first provide a brief analysis of its quality, then give a comprehensive analysis with scores for each dimension. The output must strictly follow the JSON format: ”Analysis”: …, ”Relevance”: …, ”Accuracy”: …, ”Coherence”: …, ”Clarity”: …, ”Breadth and Depth”: …, ”Reading Experience”: …. You do not need to consider whether the response meets the user’s length requirements in your evaluation. Ensure that only one integer between 1 and 5 is output for each dimension score.”’
8.3 Video-Caption Relevance Score
@@
You are an intelligent chatbot designed for evaluating the detail orientation of generative outputs for video-based question-answer pairs. ” ”Your task is to compare the predicted answer with the correct answer and determine its level of detail, considering both completeness and specificity. Here’s how you can accomplish the task:” ”——” INSTRUCTIONS: ”- Check if the predicted answer covers all major points from the video. The response should not leave out any key aspects.” ”- Evaluate whether the predicted answer includes specific details rather than just generic points. It should provide comprehensive information that is tied to specific elements of the video.” ”- Consider synonyms or paraphrases as valid matches.” ”- Provide a single evaluation score that reflects the level of detail orientation of the prediction, considering both completeness and specificity.” Please evaluate the following video-based question-answer pair: $\langle User\quad Request\rangle$ : $\langle Correct\quad Answer\rangle$ : $\langle Predicted\quad Answer\rangle$ : ”Provide your evaluation only as a detail orientation score where the detail orientation score is an integer value between 0 and 5, with 5 indicating the highest level of detail orientation. ” ”Please generate the response in the form of a Python dictionary string with keys ’score’, where its value is the detail orientation score in INTEGER, not STRING.” ”DO NOT PROVIDE ANY OTHER OUTPUT TEXT OR EXPLANATION. Only provide the Python dictionary string. ” ”For example, your response should look like this: ’score’: 4.8.”
8.4 Frame-Level Captioning

### v2

@@ 4 LongCaption-Bench: A Benchmark for Long-Caption Generation
To reliably evaluate the long-caption generation capabilities of LMMs, we constructed a LongCaption-Bench. Specifically, to prevent test videos from being included in the training data, we selected videos from the test sets of open-source long video benchmarks, with video durations ranging from 5 to 30 minutes. To reduce the difficulty of manual annotation, we first used the latest proprietary LMM, Gemini 1.5 Pro [10] (which supports long-caption generation), to generate an initial caption for all videos. We then manually reviewed each video and its corresponding initial caption, removing any abnormal captions and their associated samples. Abnormal captions included those with infinite repetition, incomplete captions (abrupt termination), and sensitive content. In the end, we retained 281 valid video-caption pairs. Subsequently, we manually reviewed the initial captions against the video content and corrected or supplemented any errors or omissions in the captions. In the appendix, we present the tool interface used for modifying and supplementing the captions. Table 1 reports the statistics of LongCaption-Bench. The test videos are primarily concentrated in the 900 to 1,200-second range, with caption lengths mostly between 500 and 1,500 words. The average video duration is 1060.4s, and the average caption length is 1161.3 words, which is significantly greater than previous video captioning benchmarks.
@@
5 LongCaptioning: A Long Video Caption Generation Model
In this section, we discuss the training details of the long-caption generation model trained on the constructed LongCaption-10K dataset, as well as the corresponding experimental results.
@@ 5.1 Model Training
We conduct training based on the latest open-source models, namely MiniCPMV2.6-8B [14]. Figure 5 shows the structure of our model. This model has only undergone pretraining and supervised fine-tuning (SFT) on image-text datasets. This allows us to eliminate the interference of other video datasets when analyzing the experimental results. Moreover, to further extend the effective context window during inference and allow the input of longer frame sequences, we introduced a visual context window extension technique [12] during training. This technique scales the rotational frequency of the position encoding embedding for visual tokens, thereby extending the model’s context window. The model was trained on a 2 A800-80G GPUs using DeepSpeed + ZeRO3 + CPU offloading [54]. We used a batch size of 1, a learning rate of 5e-5, and trained for 1 epoch.
Figure 7: Statistical results of the top 50 words in LongCaption-10K.
Figure 8: Statistical results of the top 50 words in LongCaption-Bench.
@@ 5.2.1 Quatitative Results
Evaluates the Relevance of Video and Generated Caption To ensure that the descriptions generated by the model are relevant to the video content, Table 2 presents the correlation scores between the input video and the generated descriptions. We adopted the method from Video-ChatGPT [25], using GPT-4o mini to rate the correlation between human annotations and the model-generated descriptions, with scores ranging from 1 to 5. Detailed scoring criteria and prompt templates are provided in the appendix. The experimental results indicate that our model outperforms others in terms of correlation scores across all video duration intervals. Specifically, compared to the baseline model MiniCPMV2.6-8B, our method achieved a 0.41 improvement in correlation scores. Even when compared to larger models such as Qwen2-VL-72B and GPT-4o, our model still achieved the best performance. This suggests that our model is more accurate and comprehensive in describing the input video content compared to others. Notably, MiniCPMV2.6-8B even outperformed Qwen2-VL-72b, which may be due to its training on image datasets. We found that, compared to current video captioning datasets, image captioning datasets exhibit longer caption lengths and higher quality.
Evaluates the Output Length Table 3 presents the results of the output length score of the models. We report the output length and quality score for videos of different duration intervals. For the output length score $S_{l}$, we use human annotations as the ground truth (assuming the human annotations contain all the content of the test videos) and calculate the difference between the model output length and the human annotation length using Equation 2. A higher $S_{l}$ indicates that the length of the captions generated by the model more closely matches the length of the human annotations. Our findings are as follows: 1) Open-source LMMs generally perform poorly, even those with 72 billion parameters, as they struggle to generate complete captions beyond 300 words. 2) The performance of MiniCPMV2.6-8B even surpasses that of Qwen2-VL 72B, reflecting that the output length of LMMs is not necessarily correlated with model size. 3) Notably, compared to other duration intervals, all models show an improvement in output length scores in the [1200s, 1800s) interval. After analyzing the generated outputs, we found that some models exhibit repetitive output issues when the input frame is longer, leading to an increase in output length. Due to the repetition problem, their quality scores are lower. 4) By fine-tuning the model on LongCaption-10K, our model achieved optimal performance, even surpassing larger proprietary models like GPT-4o.
Figure 6 presents the output length statistics for various models, with anomalous outputs like infinitely looping captions manually removed. Open-source models, such as MiniCPMV2.6-8B and Qwen2-VL 72B, generated captions primarily between 150 and 300 words. The proprietary GPT-4o model produced captions mostly in the 300 to 600-word range. In contrast, our model generated outputs between 300 and 700 words, with a maximum length reaching approximately 1,200 words. This demonstrates our model’s ability to produce longer, more detailed captions compared to both open-source and proprietary models.
@@ 5.2.2 Ablation Studies
We explore the effects of removing the visual context window extension and the LongCaption-10K dataset on the LongCaptioning-8B model. As shown in Table 4 and Table 5, we find that: 1) Removing the visual context window extension significantly reduces both output length and quality, with the length score $S_{l}$ dropping by 8.6, the quality score $S_{q}$ dropping by 2.1, and the relevance score $S_{r}$ dropping by 0.24. This is because the visual context window extension allows the model to input more sampled frames during the inference process. 2) Removing the clip-level captioning step during the synthesis of LongCaption-10K leads to a drop in all metrics, as relying solely on frame-level captions fails to capture dynamic changes and action continuity, resulting in less coherent captions. 3) Similarly, removing the frame-level captioning step further degrades performance, as using only clip-level captions during LongCaption-10K synthesis results in shorter captions. 4) Without the LongCaption-10K dataset ($i.e.$, using only the MiniCPMV2.6-8B backbone), the model struggles to generate long captions, leading to a sharp decline in output length. Additionally, without both the visual context window extension and the LongCaption-10K dataset, the relevance between generated captions and video content significantly deteriorates.
5.2.3 The Results of VideoMME
[…]
5.2.4 The Visualization of LongCaption-10K and LongCaption-Bench
Figures 7 and 8 display the top 50 word frequency statistics for LongCaption-10K and LongCaption-Bench, respectively. Notably, we excluded non-informative stop words (e.g., "a", "an", "the", "and", "of") to better reveal the primary themes in long-form descriptions. A significant divergence in high-frequency terms is clearly observed between LongCaption-10K and LongCaption-Bench, demonstrating notable domain discrepancy between the two datasets. This inherent distributional difference effectively mitigates potential performance inflation caused by model overfitting to specific domains during evaluation.
5.2.5 The Visualization of Long Video Captioning
Figure 9 illustrates the qualitative results of video captions generated by the comparison models (MiniCPMV2.6-8B and Qwen2-VL 72B) and our model (LongCaptioning 8B). For all models, we uniformly sample 128 frames from the video as input. It is evident that LongCaptioning 8B produces outputs with significantly greater length compared to the other models while maintaining the relevance of the generated content to the video. In contrast to MiniCPMV2.6-8B and Qwen2-VL 72B, LongCaptioning 8B generates more detailed captions for the input video.
@@ 6 Conclusions
Appendix A Training Details
@@
Table 7 reports the details of the training process. We conduct training based on the latest open-source models, namely MiniCPMV2.6-8B [14]. This model has only undergone pretraining and supervised fine-tuning (SFT) on image-text datasets. This allows us to eliminate the interference of other video datasets when analyzing the experimental results. We employ the LoRA strategy for supervised fine-tuning. Specifically, we fine-tune only a small subset of parameters in the LLM, where the $\text{LoRA}_{r}$ is set to 8 and the $\text{LoRA}_{\alpha}$ is set to 32. Moreover, to further extend the effective context window during inference and allow the input of longer frame sequences, we introduced a visual context window extension technique [12] during training. This technique scales the rotational frequency of the position encoding embedding for visual tokens, thereby extending the model’s context window. The model was trained on a 2 A800-80G GPUs using DeepSpeed + ZeRO3 + CPU offloading [54]. We used a batch size of 1, a learning rate of 5e-5, and trained for 1 epoch.
Appendix B Prompts
B.1 Test Prompt
In Figure 1, all models use the same prompt.
Task: You are an expert in understanding scene transitions based on visual features in a video. You are requested to create the descriptions for the current video sent to you, which includes multiple sequential frames. Guidelines For Video Description: - Analyze the narrative progression implied by the sequence of frames, interpreting the sequence as a whole. - Note that since these frames are extracted from a video, adjacent frames may show minimal differences. These should not be interpreted as special effects in the video. - If text appears in the frames, you must describe the text in its original language and provide an English translation in parentheses. For example: book. Additionally, explain the meaning of the text within its context. - When referring to people, use their characteristics, such as clothing, to distinguish different people. - **IMPORTANT** Please provide as many details as possible in your description, including colors, shapes, and textures of objects, actions and characteristics of humans, as well as scenes and backgrounds. Output Format: Your response should look like this: "Video Level Description": "The video begins with…, progresses by…, and concludes with…" Please give me the description of the current video.
B.2 Quality Score
@@
You are an expert in evaluating text quality. Please evaluate the quality of an AI assistant’s response to a user’s writing request. Be as strict as possible. You need to evaluate across the following six dimensions, with scores ranging from 1 to 5. The scoring criteria from 5 to 1 for each dimension are as follows: 1. Relevance: From content highly relevant and fully applicable to the user’s request to completely irrelevant or inapplicable. 2. Accuracy: From content completely accurate with no factual errors or misleading information to content with numerous errors and highly misleading. 3. Coherence: From clear structure with smooth logical connections to disorganized structure with no coherence. 4. Clarity: From clear language, rich in detail, and easy to understand to confusing expression with minimal details. 5. Breadth and Depth: From both broad and deep content with a lot of information to seriously lacking breadth and depth with minimal information. 6. Reading Experience: From excellent reading experience, engaging and easy to understand content to very poor reading experience, boring and hard to understand content. Please evaluate the quality of the following response to a user’s request according to the above requirements. $\langle User\quad Request\rangle$ : $\langle/Response\rangle$ : ”’Please evaluate the quality of the response. You must first provide a brief analysis of its quality, then give a comprehensive analysis with scores for each dimension. The output must strictly follow the JSON format: "Analysis": …, "Relevance": …, "Accuracy": …, "Coherence": …, "Clarity": …, "Breadth and Depth": …, "Reading Experience": …. You do not need to consider whether the response meets the user’s length requirements in your evaluation. Ensure that only one integer between 1 and 5 is output for each dimension score.”’
B.3 Video-Caption Relevance Score
@@
You are an intelligent chatbot designed for evaluating the detail orientation of generative outputs for video-based question-answer pairs. " "Your task is to compare the predicted answer with the correct answer and determine its level of detail, considering both completeness and specificity. Here’s how you can accomplish the task:" "——" INSTRUCTIONS: "- Check if the predicted answer covers all major points from the video. The response should not leave out any key aspects." "- Evaluate whether the predicted answer includes specific details rather than just generic points. It should provide comprehensive information that is tied to specific elements of the video." "- Consider synonyms or paraphrases as valid matches." "- Provide a single evaluation score that reflects the level of detail orientation of the prediction, considering both completeness and specificity." Please evaluate the following video-based question-answer pair: $\langle User\quad Request\rangle$ : $\langle Correct\quad Answer\rangle$ : $\langle Predicted\quad Answer\rangle$ : "Provide your evaluation only as a detail orientation score where the detail orientation score is an integer value between 0 and 5, with 5 indicating the highest level of detail orientation. " "Please generate the response in the form of a Python dictionary string with keys ’score’, where its value is the detail orientation score in INTEGER, not STRING." "DO NOT PROVIDE ANY OTHER OUTPUT TEXT OR EXPLANATION. Only provide the Python dictionary string. " "For example, your response should look like this: ’score’: 4.8."
B.4 Frame-Level Captioning